# Dynamic Integration of Task-Specific Adapters for Class Incremental Learning

### v1

@@ Conclusion
In this paper, we propose a novel framework called Dynamic Integration of task-specific Adapters (DIA), which consists of Task-Specific Adapter Integration (TSAI) and Patch-Level Model alignment. The TSAI module enhances compositionality through a patch-level adapter integration mechanism, minimizing task interference while preserving knowledge from old tasks. Patch-level model alignment maintains feature consistency and accurate decision boundaries via a patch-level distillation loss and a patch-level feature reconstruction method. 1) Our PDL preserves feature-level consistency between successive models by implementing a distillation loss based on the contribution of patch tokens to new class learning. 2) Our PFR facilitates accurate classifier alignment by reconstructing features from previous tasks that adapt to new task knowledge. We evaluate DIA on four benchmark datasets proving its efficiency and superior performance.

### v2

@@ 6 Conclusion
In this paper, we propose a novel framework called Dynamic Integration of task-specific Adapters (DIA), which consists of Task-Specific Adapter Integration (TSAI) and Patch-Level Model Alignment. First, the TSAI module enhances compositionality through a patch-level adapter integration mechanism, minimizing task interference while preserving old task knowledge. Afterward, the Patch-level Model Alignment maintains feature consistency and accurate decision boundaries via a patch-level distillation loss and a patch-level feature reconstruction method, respectively. 1) Our PDL preserves feature-level consistency between successive models by implementing a distillation loss based on the contribution of patch tokens to new class learning. 2) Our PFR facilitates accurate classifier alignment by reconstructing features from previous tasks that adapt to new task knowledge.
7 Acknowledgment
This work was supported by the National Key Research and Development Project of China (Grant 2020AAA0105600), the National Natural Science Foundation of China (Grant No. U21B2048), Shenzhen Key Technical Projects (Grant CJGJZD2022051714160501) and Natural Science Foundation of Shaanxi Province 2024JC-YBQN-0637.